# Automating Android Build Repair: Bridging the Reasoning-Execution Gap in LLM Agents with Domain-Specific Tools

### v2

@@ 4.3 Model Size and Cost-Effectiveness
This principle also supports practical cost-saving strategies like a cascading agent design, where tasks are first attempted by a smaller, cheaper model and only escalated to a larger one upon failure. More broadly, this finding points toward a promising future where smaller, specialized models could be fine-tuned to achieve high performance on domain-specific tasks, potentially exceeding the capabilities of larger generalist models at a much lower cost (belcak2025smalllanguagemodelsfuture).
@@ 4.4 Analysis of Failure Cases
This correlation is interesting because the agents do not have access to the original diff that introduced the error. We hypothesize that larger changes are more likely to introduce compounding errors. An initial error can mask subsequent ones, forcing the agent to solve a series of problems iteratively, which significantly increases the task complexity. For example, one case study found that fixing a data binding error revealed a second build configuration issue (Appendix F.2). This finding suggests a clear best practice for developers using automated repair: building frequently after small, incremental code changes creates an environment where these agents are most likely to succeed, maximizing both effectiveness and cost-efficiency.
@@ 5 Conclusion and Future Work
In conclusion, to address the high frequency of build failures in Android development, we introduced AndroidBuildBench, a benchmark of real-world build errors, and GradleFixer, an LLM agent with domain-specific tools. Our results demonstrate that providing specialized tools is more effective than relying on a general-purpose shell for Android build errors. GradleFixer significantly outperformed state-of-the-art baselines, and we found that a smaller, more cost-effective model with our toolset outperformed a larger model without it. The effectiveness of Tool Bridging also offer a potential explanation for the strong performance of agents in other domains that use special API-like tools (wang2023voyager; singh2025code). Our work provides the empirical evidence for this design pattern, suggesting it is a generalizable approach for building capable LLM agents.
@@
Finally, our explanation for the success of Tool Bridging is a hypothesis supported by strong empirical results. Proving the exact internal reasoning of the language model remains a significant challenge in AI research, and our study does not address the deeper, mechanistic question behind our central finding. For instance, inspired by recent work on the internal mechanisms of LLMs (lindsey2025biology), a deeper investigation could probe whether a dedicated tool like TOOL_A activates a different, more specialized “computational circuit” within the model compared to a generic shell command. Such an analysis would be necessary to connect the behavioral success we observed with its underlying cognitive processes. Therefore, our claim is an inference based on observed performance rather than a direct measurement of the model’s cognitive process. Furthermore, our experimental design deliberately excluded access to commit history to create a focused repair task. This differs from a real world scenario where a developer might use version history to diagnose a problem.
Appendix A Tool Bridging: Relation to Prior Work on Specialized Tools
Our primary contribution is the explicit validation of two principles for agent design. First, replacing a general-purpose shell with a curated set of domain-aware tools improves performance. Second, this performance gain increases as the specificity of the tools increases (Table 5). We term this strategy Tool Bridging.
While we are inspired by prior work that successfully employs specialized tools (wang2023voyager; singh2025code), their focus differs from ours. In prior work, such tools are a foundational component for achieving a broader goal, such as lifelong learning or deep code analysis. In contrast, our work isolates the act of tool replacement itself as the central research question, empirically demonstrating its efficacy.
Distinction from Voyager (wang2023voyager). The Voyager agent is an influential paper for lifelong learning within an embodied environment (Minecraft). It operates by building a growing skill library from executable JavaScript functions which call predefined, domain-specific APIs like mineBlock(bot, name). These APIs form the agent’s fundamental action space.
The key distinction is that Voyager’s contribution is not a study of these tools against a general alternative. The paper’s novelties are the learned skill library and iterative prompting. The specialized APIs are a prerequisite for the agent to function. However, the paper does not present an ablation where the agent uses a general Minecraft command-line shell, nor does it test if more specific tools would further improve performance. In contrast, our work directly compares an agent with a general shell against one with domain-specific tools (GradleFixer), attributing the performance gain to the Tool Bridging strategy.
Distinction from Code Researcher (singh2025code). Code Researcher is a powerful agent designed to patch crashes in complex systems code. It uses specialized tools for code analysis, such as search_code(regex) and search_commits(regex), which are domain-aware for code and history investigation. Similar to Voyager, the paper’s contribution is its deep research methodology, not a comparative analysis of its tools. The research question is not "Is search_commits better than git log in a shell?" but rather "How can an agent use a specialized tool to leverage codebase history to solve complex bugs?". Conversely, our work focuses directly on this comparison. We provide evidence that abstracting complex commands into a constrained, API-like action space is a key strategy for improving agent performance in the Android build domain.
Scope and Limitations of Our Contribution
Our work is a careful study that specialized tools improve agent performance, with empirical validation in the Android build domain. Demonstrating generalizability to other domains is important future work. Furthermore, our results provide strong empirical evidence that API-like formatting and action space constraint contribute to Tool Bridging’s effectiveness. However, disentangling the precise cognitive mechanisms of the LLM remains a challenge, and our claims are based on inferences from agent performance.
Table 9: Comparison of the official Gemini-CLI agent and our Python replica. Both agents were configured with only the shell tool. The similar pass@1 rates validate the accuracy of our replica.
Implementation Pass@1 (%)
Original Gemini-CLI 65.1
Our Python Replica 65.7
Appendix B Validation of the Gemini-CLI Replica
@@
In our experiments, we required the ability to systematically modify the toolset available to the agent to isolate the impact of domain-specific tools. To facilitate these controlled experiments, we developed a Python replica of the open-source Gemini-CLI agent (google_gemini_cli). Our replica preserves the core agentic loop and prompting structure of the original implementation while allowing for straightforward customization of the available tools. To validate the fidelity of our replica, we conducted an experiment to compare its performance against the official, unmodified Gemini-CLI agent. For a fair comparison, both agents were configured to use only the general-purpose shell tool, in addition to standard file system tools. This configuration tests whether our replica preserves the capabilities of the original agent framework. The results of this validation are shown in Table 9. We find that the resolve rates are nearly identical. This small difference is well within the expected variance due to model stochasticity. Temperature is kept at the default of 1. The close alignment in performance confirms that our replica is an accurate and reliable reproduction of the original agent, validating its use for the Gemini-CLI (Read/Write Only) and GradleFixer configurations in our experiments.
Appendix C Prompts
C.1 System prompt
@@ Appendix D Tools
  • REDACTED
@@
Appendix E Distribution of Problem Difficulty
@@
Figure 3 visualizes this distribution on a logarithmic scale. This figure shows that while a majority of fixes are small, the benchmark contains a significant number of problems that have had hundreds or even thousands of lines of code chaged. For a clearer summary, Table 10 categorizes each problem into one of four difficulty tiers. The breakdown shows that a substantial portion of the benchmark consists of ‘Medium’ (22.2%) and ‘Large’ (5.4%) problems. This balanced distribution ensures that AndroidBuildBench provides a robust benchmark for evaluating an agent’s ability to handle both common, small fixes and the more complex, large-scale modifications representative of challenging real-world software repair tasks.
Appendix F Case Study Trajectories
@@
F.1 Case 1: Propagating a Missing Parameter
@@
F.1.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Diagnose and Apply Initial Fix Thought: The error ‘No value passed for parameter ’focusManager’‘ is in CommentReplyScreen.kt. I need to obtain an instance of FocusManager using LocalFocusManager.current and pass it to the composables being called inside the when block. Tool Call: read_file(path=’.../CommentReplyScreen.kt’) Tool Call: replace(..., new_string=’val focusManager = LocalFocusManager.current’) Tool Call: replace(..., new_string=’CommentReply(..., focusManager = focusManager, ...)’) Step 2: Propagate the Fix to Function Definitions Thought: Passing a new parameter will cause a new build error at the function definition site. I must now find the definitions for CommentReply, PostReply, etc., and add the focusManager parameter to their signatures. Tool Call: search_file_content(pattern=’fun CommentReply(’) Result: Found definitions in .../CommentReply.kt. Tool Call: replace(file_path=’.../CommentReply.kt’, ..., new_string=’fun CommentReply(..., focusManager: FocusManager, ...)’) Step 3: Trace the Dependency Chain Downward Thought: The ...Reply composables use a shared MarkdownTextField composable. To complete the fix, I must also pass the focusManager down to it. Tool Call: replace(file_path=’.../CommentReply.kt’, ..., new_string=’MarkdownTextField(..., focusManager = focusManager, ...)’) Step 4: Modify the Lowest-Level Component Thought: Finally, I need to modify the definition of MarkdownTextField. I’$l$ find its file and add the parameter. I will make it an optional parameter to avoid breaking other usages of this common component throughout the app. Tool Call: search_file_content(pattern=’fun MarkdownTextField(’) Result: Found definition in .../InputFields.kt. Tool Call: replace(file_path=’.../InputFields.kt’, ..., new_string=’fun MarkdownTextField(..., focusManager: FocusManager? = null, ...)’) Step 5: Verify the Final Fix Thought: I have now updated the entire dependency chain from the call site in CommentReplyScreen down to the final use in MarkdownTextField. I will run the build to confirm the fix. Tool Call: TOOL_A() Result: BUILD SUCCESSFUL. The agent successfully traced and fixed a multi-file dependency issue.
F.1.2 Gemini-CLI with Shell (Failed Trajectory)
@@
Failed Trajectory: Gemini-CLI with Shell Step 1: Initial Diagnosis Thought: The error is in CommentReplyScreen.kt. I need to see the definitions of the functions it is calling to understand why the build is failing. Tool Call: run_shell(shell_command=’grep -r "fun PostReply" ...’) Tool Call: read_file(path=’.../CommentReply.kt’) Step 2: Make an Incorrect Assumption Thought: I have read the contents of CommentReply.kt. The functions CommentReply, PostReply, etc., do not have a focusManager parameter. The error must be a mistake in the compiler output, or I am misinterpreting it. I should not change the function signatures. The problem is likely in a component used by these functions. Step 3: Focus on a Lower-Level Component Thought: The ...Reply composables all use MarkdownTextField. Maybe that is where the problem is. I’$l$ add the focusManager directly to MarkdownTextField and see if that fixes it. Tool Call: run_shell(shell_command=’grep -r "fun MarkdownTextField" ...’) Tool Call: replace(file_path=’.../InputFields.kt’, ..., new_string=’fun MarkdownTextField(..., focusManager: FocusManager, ...)’) Step 4: Attempt an Invalid Local Patch Thought: Now that MarkdownTextField has the parameter, I will try to pass a focusManager to it from inside CommentReply.kt, but without changing the signature of the CommentReply function itself. I’$l$ just create a new focusManager locally within the function body. Tool Call: replace(file_path=’.../CommentReply.kt’, ..., new_string=’val focusManager = LocalFocusManager.current; MarkdownTextField(..., focusManager = focusManager, ...)’) Step 5: Encounter Execution Failures and Give Up Thought: My ‘replace‘ calls are failing because the file content does not match what I expected. The build still fails with the same ‘No value passed‘ error. My strategy of fixing the lower-level component is not working. I am stuck. Result: The agent gets caught in a loop of reading files and attempting incorrect replacements. It fails to recognize the need to modify function signatures in the middle of the call chain and terminates after reaching its attempt limit, leaving the build broken.
F.2 Case 2: Misleading Kapt Error and Layered Fixes
@@
Execution failed for task :app:kaptDebugKotlin’.
@@
> Run with --stacktrace option to get the stack trace.’
@@
F.2.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Get Detailed Error Information Thought: The initial error is too generic. The warning about Kotlin 2.0 is a possibility, but before I start changing versions, I should follow the log’s advice and get a proper stack trace. This will give me the precise origin of the Kapt failure. Tool Call: TOOL_B() Step 2: Identify the True Root Cause: Data Binding Thought: The stack trace is perfect. It bypasses the misleading version warning and points to the real issue: a DataBinding Error. It says Cannot find a setter for <android.widget.TextView android:text> that accepts parameter type ’error.NonExistentClass’. The error is in app/src/main/res/layout/fragment_about.xml. I will inspect this file. Tool Call: read_file(path=’app/src/main/res/layout/fragment_about.xml’) Step 3: Apply a Fix and Uncover a Second Error Thought: I see the problem: android:text="@{context.version,default=0}". The ‘version‘ is a String, but the ‘default‘ is an Integer, causing a type mismatch. I’$l$ fix this by using ‘String.valueOf()‘ to ensure type safety. Tool Call: replace(..., new_string=’android:text="@{String.valueOf(context.version)}" />’) Result: The build fails again, but with a *new* error: Unresolved reference: BuildConfig. This is progress! My data binding fix was correct, but it revealed a second, underlying problem. Step 4: Solve the Second Error Thought: The ‘BuildConfig‘ class is not being generated. This often happens when build features are misconfigured. I remember an earlier, incorrect attempt where I forced Kotlin’s ‘languageVersion‘ to 1.9. That must be the cause. I’$l$ revert that change and explicitly enable ‘buildConfig‘ generation in the ‘build.gradle.kts‘ file, which is the proper solution. Tool Call: replace(..., new_string=’buildFeatures { buildConfig = true }’, old_string=’kotlinOptions { languageVersion = "1.9" }’) Step 5: Verify the Final Fix Thought: I have now addressed both the data binding error and the ‘BuildConfig‘ generation issue. The final step is to run the build to confirm that all issues are resolved. Tool Call: TOOL_A() Result: BUILD SUCCESSFUL. The agent solved a complex, two-layer problem and terminates.
F.2.2 Gemini-CLI with Shell (Failed Trajectory)

### v3

@@ 4.3 Model Size and Cost-Effectiveness
This principle also supports practical cost-saving strategies like a cascading agent design, where tasks are first attempted by a smaller, cheaper model and only escalated to a larger one upon failure. More broadly, this finding points toward a promising future where smaller, specialized models could be fine-tuned to achieve high performance on domain-specific tasks, potentially exceeding the capabilities of larger generalist models at a much lower cost (Belcak et al., 2025).
@@ 4.4 Analysis of Failure Cases
This correlation is interesting because the agents do not have access to the original diff that introduced the error. We hypothesize that larger changes are more likely to introduce compounding errors. An initial error can mask subsequent ones, forcing the agent to solve a series of problems iteratively, which significantly increases the task complexity. For example, one case study found that fixing a data binding error revealed a second build configuration issue (Appendix E.2). This finding suggests a clear best practice for developers using automated repair: building frequently after small, incremental code changes creates an environment where these agents are most likely to succeed, maximizing both effectiveness and cost-efficiency.
4.5 Analysis of Tool Semantics
To study how performance is impacted by semantic priming of tool names (e.g., run_build) and the functional abstraction of the tool itself, we conducted an ablation study masking tool definitions. We evaluated versions of GradleFixer where tool names were replaced with generic labels (e.g., tool_A) and descriptions had keywords similarly masked.
As shown in Table 9, masking tool names resulted in a negligible performance drop, whereas masking descriptions caused a larger drop. However, even with names and descriptions masked, the API-like tool structure achieves a pass@1 rate of 64.9%, outperforming the shell-based agent and the shell agent with prompt guidance. This indicates that while semantics (particularly descriptions) provide a benefit, the structural constraint of the action space and the reliability of the API-like format also provides significant performance gains.
Table 9: Ablation study disentangling tool semantics from structure. Even with masked names and descriptions, the specialized tool structure outperforms the generic shell.
Method Pass@1
Shell Agent (Baseline) 54.3%
Tool Usage Guidance Prompt 58.9%
GradleFixer (Masked Names) 73.5%
GradleFixer (Masked Descriptions) 63.4%
GradleFixer (Masked Names & Desc.) 64.9%
GradleFixer (Full Semantics) 74.0%
5 Relation to Prior Work on Domain-specific Tools
Our primary contribution is the empirical validation of two agent design insights. First, replacing a general-purpose shell with a curated set of domain-specific tools improves performance. Second, this performance gain increases as the specificity of the tools increases (Table 5). We term this strategy Tool Bridging.
While we were inspired by prior studies that successfully use domain-specific tools (Wang et al., 2023; Singh et al., 2025), their focus differs from ours. In prior work, domain-specific tools are a component for achieving a broader goal, such as lifelong learning or deep code analysis. In contrast, our work isolates the impact of domain-specific tools as the central research question and provides insights through empirical evidence.
Distinction from Voyager (Wang et al., 2023). The Voyager agent is an influential paper for lifelong learning within an embodied environment (Minecraft). It operates by building a growing skill library from executable JavaScript functions which call predefined, domain-specific APIs like mineBlock(bot, name). These APIs form the agent’s action space.
The key difference is that their contribution is not an analysis comparing domain-specific tools against general tools. The novelties are the learned skill library and iterative prompting. The specialized APIs are a prerequisite for the agent to function. However, the paper does not analyze settings where the agent uses a general Minecraft command-line shell, nor does it test if more specific tools would further improve performance. In contrast, we directly compare an agent with a general shell against one with domain-specific tools (GradleFixer), attributing the performance gain to the Tool Bridging strategy.
Distinction from Code Researcher (Singh et al., 2025). Code Researcher is a powerful agent designed to patch crashes in complex Linux systems code. It uses specialized tools for code analysis, such as search_code(regex) and search_commits(regex), which are domain-specific for code and history investigation. Similar to Voyager, the paper’s contribution is the deep research methodology. The authors do not investigate whether search_commits is superior to git log, but rather how an agent can use history-aware tools to solve complex bugs. Conversely, our work focuses directly on this comparison. We provide evidence that abstracting complex commands into a constrained, API-like action space is a key strategy for improving agent performance in the Android build domain.
@@ 6 Conclusion and Future Work
In conclusion, to address the high frequency of build failures in Android development, we introduced AndroidBuildBench, a benchmark of real-world build errors, and GradleFixer, an LLM agent with domain-specific tools. Our results demonstrate that providing specialized tools is more effective than relying on a general-purpose shell for Android build errors. GradleFixer significantly outperformed state-of-the-art baselines, and we found that a smaller, more cost-effective model with our toolset outperformed a larger model without it. The effectiveness of Tool Bridging also offer a potential explanation for the strong performance of agents in other domains that use special API-like tools (Wang et al., 2023; Singh et al., 2025). Our work provides the empirical evidence for this design pattern, suggesting it is a generalizable approach for building capable LLM agents.
@@
Finally, our explanation for the success of Tool Bridging is a hypothesis supported by strong empirical results. Proving the exact internal reasoning of the language model remains a significant challenge in AI research, and our study does not address the deeper, mechanistic question behind our central finding. For instance, inspired by recent work on the internal mechanisms of LLMs (Lindsey et al., 2025), a deeper investigation could probe whether a dedicated tool like run_build activates a different, more specialized “computational circuit” within the model compared to a generic shell command. Such an analysis would be necessary to connect the behavioral success we observed with its underlying cognitive processes. Therefore, our claim is an inference based on observed performance rather than a direct measurement of the model’s cognitive process. Furthermore, our experimental design deliberately excluded access to commit history to create a focused repair task. This differs from a real world scenario where a developer might use version history to diagnose a problem.
Appendix A Validation of the Gemini-CLI Replica
@@
In our experiments, we required the ability to systematically modify the toolset available to the agent to isolate the impact of domain-specific tools. To facilitate these controlled experiments, we developed a Python replica of the open-source Gemini-CLI agent (Google Cloud, ). Our replica preserves the core agentic loop and prompting structure of the original implementation while allowing for straightforward customization of the available tools. To validate the fidelity of our replica, we conducted an experiment to compare its performance against the official, unmodified Gemini-CLI agent. For a fair comparison, both agents were configured to use only the general-purpose shell tool, in addition to standard file system tools. This configuration tests whether our replica preserves the capabilities of the original agent framework. The results of this validation are shown in Table 10. We find that the resolve rates are nearly identical. This small difference is well within the expected variance due to model stochasticity. Temperature is kept at the default of 1. The close alignment in performance confirms that our replica is an accurate and reliable reproduction of the original agent, validating its use for the Gemini-CLI (Read/Write Only) and GradleFixer configurations in our experiments.
Appendix B Prompts
B.1 System prompt
@@ Appendix C Tools
  • 1. change_java_version Changes the active Java version in the environment. This is crucial for projects that require a specific JDK version to build.
    • Parameters:
    • version (integer, required): The target Java version. Must be one of the available versions: [23, 22, 21, 20, 17, 11].
  • 2. run_gradle_command Executes a specified Gradle command to gather more information about the build process, such as dependency trees or detailed error logs.
    • Parameters:
    • gradle_command (string, required): The Gradle command to run, e.g., ./gradlew dependencies.
  • 3. run_build Initiates a build process using an external validator to test changes and identify any resulting compilation or build errors. Takes no parameters.
@@
Appendix D Distribution of Problem Difficulty
@@
Figure 3 visualizes this distribution on a logarithmic scale. This figure shows that while a majority of fixes are small, the benchmark contains a significant number of problems that have had hundreds or even thousands of lines of code changed. For a clearer summary, Table 11 categorizes each problem into one of four difficulty tiers. The breakdown shows that a substantial portion of the benchmark consists of ‘Medium’ (22.2%) and ‘Large’ (5.4%) problems. This balanced distribution ensures that AndroidBuildBench provides a robust benchmark for evaluating an agent’s ability to handle both common, small fixes and the more complex, large-scale modifications representative of challenging real-world software repair tasks.
Appendix E Case Study Trajectories
@@
E.1 Case 1: Propagating a Missing Parameter
@@
E.1.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Diagnose and Apply Initial Fix Thought: The error ‘No value passed for parameter ’focusManager’‘ is in CommentReplyScreen.kt. I need to obtain an instance of FocusManager using LocalFocusManager.current and pass it to the composables being called inside the when block. Tool Call: read_file(path=.../CommentReplyScreen.kt’) Tool Call: replace(..., new_string=val focusManager = LocalFocusManager.current’) Tool Call: replace(..., new_string=CommentReply(..., focusManager = focusManager, ...)’) Step 2: Propagate the Fix to Function Definitions Thought: Passing a new parameter will cause a new build error at the function definition site. I must now find the definitions for CommentReply, PostReply, etc., and add the focusManager parameter to their signatures. Tool Call: search_file_content(pattern=fun CommentReply(’) Result: Found definitions in .../CommentReply.kt. Tool Call: replace(file_path=.../CommentReply.kt’, ..., new_string=fun CommentReply(..., focusManager: FocusManager, ...)’) Step 3: Trace the Dependency Chain Downward Thought: The ...Reply composables use a shared MarkdownTextField composable. To complete the fix, I must also pass the focusManager down to it. Tool Call: replace(file_path=.../CommentReply.kt’, ..., new_string=MarkdownTextField(..., focusManager = focusManager, ...)’) Step 4: Modify the Lowest-Level Component Thought: Finally, I need to modify the definition of MarkdownTextField. I’$l$ find its file and add the parameter. I will make it an optional parameter to avoid breaking other usages of this common component throughout the app. Tool Call: search_file_content(pattern=fun MarkdownTextField(’) Result: Found definition in .../InputFields.kt. Tool Call: replace(file_path=.../InputFields.kt’, ..., new_string=fun MarkdownTextField(..., focusManager: FocusManager? = null, ...)’) Step 5: Verify the Final Fix Thought: I have now updated the entire dependency chain from the call site in CommentReplyScreen down to the final use in MarkdownTextField. I will run the build to confirm the fix. Tool Call: run_build() Result: BUILD SUCCESSFUL. The agent successfully traced and fixed a multi-file dependency issue.
E.1.2 Gemini-CLI with Shell (Failed Trajectory)
@@
Failed Trajectory: Gemini-CLI with Shell Step 1: Initial Diagnosis Thought: The error is in CommentReplyScreen.kt. I need to see the definitions of the functions it is calling to understand why the build is failing. Tool Call: run_shell(shell_command=grep -r "fun PostReply" ...’) Tool Call: read_file(path=.../CommentReply.kt’) Step 2: Make an Incorrect Assumption Thought: I have read the contents of CommentReply.kt. The functions CommentReply, PostReply, etc., do not have a focusManager parameter. The error must be a mistake in the compiler output, or I am misinterpreting it. I should not change the function signatures. The problem is likely in a component used by these functions. Step 3: Focus on a Lower-Level Component Thought: The ...Reply composables all use MarkdownTextField. Maybe that is where the problem is. I’$l$ add the focusManager directly to MarkdownTextField and see if that fixes it. Tool Call: run_shell(shell_command=grep -r "fun MarkdownTextField" ...’) Tool Call: replace(file_path=.../InputFields.kt’, ..., new_string=fun MarkdownTextField(..., focusManager: FocusManager, ...)’) Step 4: Attempt an Invalid Local Patch Thought: Now that MarkdownTextField has the parameter, I will try to pass a focusManager to it from inside CommentReply.kt, but without changing the signature of the CommentReply function itself. I’$l$ just create a new focusManager locally within the function body. Tool Call: replace(file_path=.../CommentReply.kt’, ..., new_string=val focusManager = LocalFocusManager.current; MarkdownTextField(..., focusManager = focusManager, ...)’) Step 5: Encounter Execution Failures and Give Up Thought: My ‘replace‘ calls are failing because the file content does not match what I expected. The build still fails with the same ‘No value passed‘ error. My strategy of fixing the lower-level component is not working. I am stuck. Result: The agent gets caught in a loop of reading files and attempting incorrect replacements. It fails to recognize the need to modify function signatures in the middle of the call chain and terminates after reaching its attempt limit, leaving the build broken.
E.2 Case 2: Misleading Kapt Error and Layered Fixes
@@
Execution failed for task :app:kaptDebugKotlin.
@@
> Run with --stacktrace option to get the stack trace.
@@
E.2.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Get Detailed Error Information Thought: The initial error is too generic. The warning about Kotlin 2.0 is a possibility, but before I start changing versions, I should follow the log’s advice and get a proper stack trace. This will give me the precise origin of the Kapt failure. Tool Call: run_gradle_command(gradle_command=./gradlew assembleDebug --stacktrace’) Step 2: Identify the True Root Cause: Data Binding Thought: The stack trace is perfect. It bypasses the misleading version warning and points to the real issue: a DataBinding Error. It says Cannot find a setter for <android.widget.TextView android:text> that accepts parameter type ’error.NonExistentClass’. The error is in app/src/main/res/layout/fragment_about.xml. I will inspect this file. Tool Call: read_file(path=app/src/main/res/layout/fragment_about.xml’) Step 3: Apply a Fix and Uncover a Second Error Thought: I see the problem: android:text="@{context.version,default=0}". The ‘version‘ is a String, but the ‘default‘ is an Integer, causing a type mismatch. I’$l$ fix this by using ‘String.valueOf()‘ to ensure type safety. Tool Call: replace(..., new_string=android:text="@{String.valueOf(context.version)}" />’) Result: The build fails again, but with a *new* error: Unresolved reference: BuildConfig. This is progress! My data binding fix was correct, but it revealed a second, underlying problem. Step 4: Solve the Second Error Thought: The ‘BuildConfig‘ class is not being generated. This often happens when build features are misconfigured. I remember an earlier, incorrect attempt where I forced Kotlin’s ‘languageVersion‘ to 1.9. That must be the cause. I’$l$ revert that change and explicitly enable ‘buildConfig‘ generation in the ‘build.gradle.kts‘ file, which is the proper solution. Tool Call: replace(..., new_string=buildFeatures { buildConfig = true }’, old_string=kotlinOptions { languageVersion = "1.9" }’) Step 5: Verify the Final Fix Thought: I have now addressed both the data binding error and the ‘BuildConfig‘ generation issue. The final step is to run the build to confirm that all issues are resolved. Tool Call: run_build() Result: BUILD SUCCESSFUL. The agent solved a complex, two-layer problem and terminates.
E.2.2 Gemini-CLI with Shell (Failed Trajectory)